write the rule for thr nth term of the arithmetic sequence.
1) 5,,14,23,32,41
2)d=-3;a6=30
3)-1,-4,-7,-10,-13... 4) d=5;a5=40

Answer by Edwin McCravy(20054)

The formula to learn for the nth term of arithmetic
sequences is:

an = a1 + (n-1)·d

We determine d by subtracting any term from the next term:

5,14,23,32,41,.....
a2 - a1 = 14 - 5 = 9
a3 - a2 = 23 - 14 = 9
a4 - a3 = 32 - 23 = 9
a5 - a4 = 41 - 32 = 9

So d = 9

an = a1 + (n-1)·d
an = 5 + (n-1)·9
an = 5 + 9(n-1)
an = 5 + 9n - 9
an = 9n - 4

---------------------------------------

d=-3;a6=30
an = a1 + (n-1)·d


Substitute n=6

a6 = a1 + (6-1)·(-3)
30 = a1 + (5)(-3)
30 = a1 - 15
45 = a1

an = a1 + (n-1)·(-3)
an = 45 + (-3)(n-1)
an = 45 - 3(n-1)
an = 45 - 3n + 3
an = 48 - 3n

----------------------------------------

-1,-4,-7,-10,-13,....
a2 - a1 = -4 - (-1) = -4 + 1 = -3
a3 - a2 = -7 - (-4) = -7 + 4 = -3
a4 - a3 = -10 - (-7) = -10 + 7 = -3 
a5 - a4 = -13 - (-10) = -13 + 10 = -3

So d = -3

an = a1 + (n-1)·d
an = -1 + (n-1)·(-3)
an = -1 + (-3)(n-1)
an = -1 - 3(n-1)
an = -1 - 3n + 3
an = 2 - 3n 

---------------------------------------      

d=5;a5=40

an = a1 + (n-1)·d


Substitute n=5

a5 = a1 + (5-1)·(5)
40 = a1 + (4)(5)
40 = a1 + 20
20 = a1

an = a1 + (n-1)·(5)
an = 20 + 5(n-1)
an = 20 + 5n - 5
an = 15 + 5n

Edwin